OpenStudy (anonymous):

You should have a minimum of 7 lines of work. 5(2x - 7) = 33 - 3(6x + 4) Next, type out in words the steps you took to solve the equation on the same document.

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

Distribute into the parenthesis. 10x-35=33-(18x+12) Distribute the negative sign 10x-35=33-18x-12 combine like terms 10x-35=21-18x put the x's in the same side 10x+18x=35+21 combine like terms 28x=56 solve for x x=2 check answer

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Don't give them the answer, help them figure it out. Giving out the answer is against the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (farmdawgnation):

Wired is correct, @Raghav77. Only providing the answer is against the CoC. Please don't do it again. Thanks. :)

OpenStudy (jiteshmeghwal9):

Distribute 5 btw 2x & -7 & 3 btw 6x & 4. 10x-35=33-18x-12 10x+18x=21+35 28x=56 x=56/28=2.

OpenStudy (anonymous):

@jiteshmeghwal9 that was already posted by @ArcticStroke

OpenStudy (jiteshmeghwal9):

Is there any problem if i have rewwritten it

OpenStudy (anonymous):

Not really, but why repost it?